# MMSD3.0: A Multi-Image Benchmark for Real-World Multimodal Sarcasm Detection

### v1

@@ 5.11 Error Analysis
Figure 7 illustrates representative failure cases revealing key challenges in multi-image sarcasm detection. In (a) and (b), the model overlooks abstract or culture-dependent incongruity. Example (a) involves metaphorical exaggeration linking visually unrelated objects through shared social context, while (b) depends on cultural knowledge and meme-level humor involving character references. In (c) and (d), subtle linguistic pragmatics—phrases like “may work,” “such a shame,” or “too overblown”—carry mixed emotional tones that confuse the sentiment classifier. These cases show that sarcasm comprehension goes beyond multimodal fusion, requiring commonsense reasoning, contextual grounding, and sensitivity to pragmatic intent.

### v2

@@ 5.11 Error Analysis
Figure 7 illustrates representative failure cases revealing key challenges in multi-image sarcasm detection. In (a) and (b), the model overlooks abstract or culture-dependent incongruity. Example (a) fails mainly because the text is too short, providing no contextual signal for interpreting the visual metaphor. The sarcastic intent relies on exaggerated visual analogy across images, which the model misses without textual grounding. In (b), the model fails to recognize sarcasm connected to meme-style humor and football culture, where the reference to “doomwatch” depends on shared cultural knowledge that the model cannot infer from surface text.
In (c) and (d), subtle linguistic pragmatics—phrases like “may work,” “such a shame,” or “too overblown”—carry mixed emotional tones that confuse the sentiment classifier. These ambiguous expressions blend positive and negative cues, making it difficult for the model to correctly detect sarcasm. Overall, these cases indicate that beyond multimodal fusion, sarcasm detection requires commonsense reasoning, contextual grounding, and deeper sensitivity to pragmatic and cultural nuances.